A roller of mass 300kg and of radius 50cm is lying on horizontal floor is resting against a step of height 20cm. The minimum horizontal force to be applied on the roller passing through its centre to turn the roller on the step.

Grade:12

1 Answers

Quasim Khan
7 years ago
mgh=1/2mv2+1/2Iw2
mgh=1/2mv2+1/2mr2v2/r2
mgh=mv2
v=root gh
v=1.414 m/s
again
W=F.v
1/2mv2+1/2mr2v2/r2=F.v
mv2=F.v
F=m.v
F=300*1.414
F=424.2 N